OpenStudy (anonymous):

mrs. morton has a special reward system for her class. when all her students behave well, she rewards them by putting 3 marbles in a jar, when the jar has 100 marbles or more the class has a party. right now, the jar contains 24 marbles. let r represent the number of additional times the class is rewarded. write an inequality to determine how many more times the class needs to be rewarded in order to earn a party. _______ what is the minimum whole number of times they need to be rewarded so they can have a party? _______ times.

OpenStudy (anonymous):

if they are rewarded \(r\) times they will get \(3r\) marbles they larady have \(24\) so you want \[24+3r\geq 100\]

OpenStudy (anonymous):

*already

OpenStudy (anonymous):

solve that for \(r\) by subtracting \(24\) and then dividing by \(3\)

OpenStudy (anonymous):

i am so confused this question is confusing nd so r tha answers

OpenStudy (anonymous):

nvm i get it

OpenStudy (anonymous):

\[24+3r\geq 100\] \[3r\geq 76\]\[r\geq 76\div 3\]

OpenStudy (anonymous):

thanks